OpenStudy (anonymous):

write the equation of a line that is parallel to 4x=3y=7 and also passes through the point (-5,2) (PLEASEEE HELP MEEEE!) i been sittin here for an hour and a half ;_;

OpenStudy (anonymous):

You mean to write 4x+3y=7 is that correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Theorem: the slope of ax + by = c is -a/b Thus slope is -4/3 Now use point-slope formula y-y1=m(x-x1) y-2=(-4/3)(x-(-5)) y=(-4/3)x -20/3 +2 y=(-4/3)x - 14/3

OpenStudy (anonymous):

THANK YOU!!!!!

OpenStudy (anonymous):

ur welcome :})